OpenStudy (anonymous):

the equations are written in point-slope form, which equation represents a line through points (-8,3)and (-2,-3) a. y-3=1(x+8) b. y-3=3(x+8) c. y-3=-3 (x+8) d. y-3=-1(x+8) suppose you and some friends are going to the movies. the amount of money you spend for tickets varies directly with the number of people buying ticket , if you spend $48 for 6 tickets. what is the constant of varitoin? a. 48 b. 8 c. 6 d. 1

jigglypuff314 (jigglypuff314):

first find the slope between the two points (-8,3) and (-2,-3) can you do that?

OpenStudy (anonymous):

no im bad a slopes

OpenStudy (anonymous):

at

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

(-8,3) and (-2,-3) Point slope form M=y2-y1/x2-x1=-3-3/-2+8=-9/6=-3/2 y-y1=m(x-x1) y-3=-3/2(x+8)

jigglypuff314 (jigglypuff314):

\[\frac{ -3-3 }{ -2-(-8) } =~?\]

OpenStudy (anonymous):

2. y=kx

jigglypuff314 (jigglypuff314):

@corderoalize -3-3 is not -9

OpenStudy (anonymous):

Constant variation would be 8

OpenStudy (anonymous):

Sorry it = -6

OpenStudy (anonymous):

Sorry the slope would be -1

jigglypuff314 (jigglypuff314):

let's let @kate_bell do parts of it herself @corderoalize to make sure she understands the process

OpenStudy (anonymous):

1. y-y1=m(x-x1) y-3=-1(x+8) 2. y=kx 48=k(6) 8=k

OpenStudy (anonymous):

They ask for help so we can give them the answers and help them through it so they can learn the process because your way is confusing

OpenStudy (anonymous):

1. D 2. B

jigglypuff314 (jigglypuff314):

*sigh* @corderoalize Please try to avoid giving such direct answers, it is against OpenStudy's code of conduct http://openstudy.com/code-of-conduct This site has a policy for "Give Help, Not Answers" violations of such policies could potentially result in an account suspension or ban. So, please keep that in mind from now on :)

OpenStudy (anonymous):

Ok but how many times have you repeated that because everyone does it

jigglypuff314 (jigglypuff314):

ahaha yep, so I have the copy-paste version xD @corderoalize

OpenStudy (anonymous):

i got it yall